OpenStudy (anonymous):

prove using identities that ((sec^2)x)/(((sec^2)x)-1)=(csc^2)x

OpenStudy (bahrom7893):

Okay, so let me try this: did you mean: \[Sec^2(x)/(Sec^2x - 1) = Csc^2(x)\]

OpenStudy (anonymous):

yea. sorry i don't know how get it like that on my computer

OpenStudy (bahrom7893):

okay so now workin on it

OpenStudy (anonymous):

thank you so much

OpenStudy (bahrom7893):

There are a few Sec and Csc identities, but I never remember them so im just gonna convert em to Sin and Cos: Sec(x) = 1/Cos(x) Csc(x) = 1/Sin(x)

OpenStudy (bahrom7893):

\[[1/Cos^2x] / [ (1/Cos^2x)-1 ] = 1/Sin^2x\]

OpenStudy (bahrom7893):

Now let's simplify this part: \[(1/Cos^2x) - 1\] = \[(1-Cos^2x)/Cos^2x\]

OpenStudy (anonymous):

wait. how did you do that?

OpenStudy (bahrom7893):

find the common denominator

OpenStudy (bahrom7893):

1 = Cos^2 / Cos^2... get it?

OpenStudy (anonymous):

oh . yea, sorry. i got it

OpenStudy (bahrom7893):

k

OpenStudy (bahrom7893):

So now you have to divide 1/Cos^2x by (1-Cos^2x)/Cos^2x which is the same as multiplying by its inverse

OpenStudy (bahrom7893):

\[[1/Cos^2x] * [Cos^2x/(1-Cos^2x)] = 1/Sin^2x\]

OpenStudy (bahrom7893):

Cancel out cos^2s

OpenStudy (bahrom7893):

\[1/(1-Cos^2x) = 1/Sin^2x\]

OpenStudy (bahrom7893):

Cross multiply:

OpenStudy (anonymous):

wait. can't you just say 1 / (sin^2x) =csc^2x

OpenStudy (anonymous):

and then be done? or is that skipping a step?

OpenStudy (bahrom7893):

yeah, but i prefer keeping everything in sin and cosine

OpenStudy (bahrom7893):

bcuz its easy to simplify

OpenStudy (bahrom7893):

\[Sin^2x = 1 - Cos^2x\]

OpenStudy (bahrom7893):

See that's proved. Sin^2x + Cos^2x = 1!!!!

OpenStudy (bahrom7893):

get it?

OpenStudy (anonymous):

yes. thank you so much!

OpenStudy (bahrom7893):

k off to helpin someone else.. any questions: email me bahrom.cfo@gmail.com and if u liked my explanation, bcum a fan YAY!

OpenStudy (anonymous):

deff will. that was so helpful. thanks again!